OpenStudy (anonymous):

At the grocery store, Lindsay bought several pounds of strawberries. The strawberries cost $2.25 per pound. After she paid the cashier, she had $8.75 remaining. Write an expression with one variable to represent the amount of money Lindsay had before she bought the strawberries. Then, use your expression to find how much money Lindsay started with before buying 4 pounds of strawberries. Medal and fan for answer.

OpenStudy (anonymous):

@Nnesha @Loser66 @phi @welshfella

OpenStudy (anonymous):

@Skielerlucas04

OpenStudy (anonymous):

idk sorry :/

OpenStudy (anonymous):

It's ok.

OpenStudy (welshfella):

y = original amount of money x = number of pounds of berries bought y - 2.25x = 8.75 y = 2.25x + 8.75

OpenStudy (welshfella):

if she bought 4 pounds then x = 4 so original amt money = y = 2.25*4 + 8.75 = $17.75

OpenStudy (anonymous):

ty

OpenStudy (welshfella):

yw any time